OpenStudy (anonymous):

Evaluate . 7C6 A. 9 B. 1 C. 5,040 D. 7 can someone explain this to me?

Parth (parthkohli):

In general: \( \color{Black}{\Rightarrow nCr= \Large {n! \over (n - r)! \times r!} }\) ! means 'factorial'.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 7C6 = \Large {7! \over (7 - 6)! \times 6!} }\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow n! = n \times (n - 1) \times (n - 2) ......\times1 }\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow 7! = 7 \times 6 \times 5 \times 4\times 3\times 2\times 1 }\)

OpenStudy (anonymous):

so the answer would be what?

OpenStudy (anonymous):

5,040 is the answer